Redraw everything using Alice’s frame of reference.
Once again, we and our friends, Alice and Bob, are inertial observers who
all meet at a single event. At this event, our clock, Alice’s clock, and Bob’s
clock all read zero and a firecracker explodes. Alice moves to our right at
c/2 and Bob moves to our left at c/2.
1
Expert's answer
2019-12-02T10:13:06-0500

in our reference frame


In the Alice reference frame



Velocity of Bob in the Alice reference frame


"v_B'=\frac{v_A+v_B}{1+\frac{v_Av_B}{c^2}}=\frac{2}{3}c"